PIP-II report for May 18, 2018

Highlights:

  • SSR1 dressed cavity S103 welding completed at BARC.
  • HWR integrated cold test with high power coupler reached 5 kW in CW mode.
  • Two SSR1 cavities with high power couplers are ready for string assembly.
  • Cryogenics team visited ESS to discuss cryoplant and cryogenics distribution system.

 

ESH&Q (Teri Dykhuis & John Anderson):

  • Ongoing communication with the Illinois Department of Natural Resources (IDNR) and United States Fish and Wildlife Service (USFWS) concerning potential impact of LBNF and PIP-II to the habitat of any endangered species that have been seen on site.
  • Developing geographic information system maps for these two agencies.
  • EA progress is ongoing.

 

Project Management (Marc Kaducak):

  • Worked with IPT on a list of items to focus on for CD-2. Some items include a design maturity approach, in-kind risk mitigation, KPPs, procurement planning, AIP/GPP interfaces, and resource management across the lab.
  • Ramped up efforts on the resource loaded schedule. Safety system section is complete and magnets/power supplies are under review. Next areas of focus include cryomodules and conventional facilities.

 

Technical Integration (Arkadiy Klebaner):

  • System managers have been engaging in interface discussions. Interface development is nearing 50% completion.
  • Continue work on the consolidated Preliminary Design Review plan.
  • Instituted Design Change Request procedure.  Initiated the procedure at HWR request.
  • Design Review Report and Review Charge Templates released and are being reviewed.
  • Developing Functional Requirements Specification (FRS) guideline document along with a template.

International Coordination (Steve Holmes & Lia Merminga):

  • DAE Labs:
    • SSR1 dressed cavity S103 welding completed at BARC

 

SRF & Cryo Systems (Allan Rowe & Genfa Wu):

HWR

  • HWR CM Production
    • FNAL and ANL staff are presently reviewing the HWR CM assembly and testing plans and procedures with the goal of optimizing the CM delivery date to Fermilab.
  • HWR Integrated Cold Test with Power Coupler
    • Integrated Cold Test (@4K) is ongoing at ANL on HWR #5 with power coupler. The input RF power reached the 5 kW level and ran in CW mode for ~30 min, when the 10 kW circulator failed. The root cause investigation is ongoing. During the 5 kW period the max observed temp rises were: +3K @5K flange, +5K @80K flange and +15K @300K flange. There is no evidence of multipacting present.
    • The RF system at ANL that supports HWR coupler and cavity testing is being upgraded with assistance from FNAL personnel to utilize a higher power circulator and improved interlocks.
    • The coupler/cavity production testing protocol is being reviewed in order to ensure that the test plan is efficient, addresses the appropriate performance concerns, and minimizes risk to components.

SSR1

  • SSR1 Cavities, coupler and tuners
    • S107 test with high power coupler is in progress.
  • SSR1 Cryomodule
    • Two cavities with high power coupler are ready for string assembly. S1H-NR-112 and S1H-NR-113.

SSR2

  • A meeting with Orsay colleagues is scheduled this Thursday to kick off collaboration on SSR2.

HB650

  • Package for procurement of Nb for prototype cavities (qty 2) being prepared. Requisition to be entered in the near future.
  • 5-cell beta=0.90 cavity B9A-AES-010 is selected for the first cavity dressing exercise.
  • Beta=0.61 cavity mechanical simulations with stiffener ring radius of 90 mm performed by VECC. Initial results indicate the cavity will not be within specifications. Further clarification and evaluation is being performed.

Cryogenics Plant

  • RLS development for the Cryoplant is in progress.
  • Collaboration visit to ESS on Tuesday of this week. The purpose of the visit is to discuss the ESS Cryoplant and supporting systems which is similar to the proposed PIP-II Cryoplant and supporting systems.

Cryogenics Distribution System

  • RLS development for the CDS is in progress.
  • Collaboration visit to ESS on Tuesday of this week. The purpose of the visit is to discuss the ESS CDS which is similar to the proposed PIP-II CDS.

 

Accelerator Systems (Paul Derwent):

High Power RF

  • The 650 MHz TRS has been approved by checkers on Teamcenter and is now out for approval among system interface designers.  We are now converging with RRCAT colleagues on the testing and verification procedures.
  • We have been supporting HWR coupler testing.  The 7kW amplifier and circulator were commissioned at ANL.  However, in the course of testing, a fault damaged the circulator.  Next step is to replace and modify the distribution on Monday.
  • The design of the RF distribution for the 650 MHz STC modifications is nearly complete.  We just need to design the transition into the cavity input coupler.
  • The first draft of the ICD is 85% complete.

Instrumentation

  • Below is a plot of our first wire scanner measurements at PIP2IT. The black trace (device P:M61WVC) is vertical profile, the yellow trace (device P:M61WHC) is horizontal profile and the blue trace (device P:MX1SPC) is the beam current. There is a small drop in beam current (~3%) during a scan as the wire interacts with the beam.

First PIP2IT wire scanner measurements. The black trace is the vertical profile, the yellow trace the horizontal profile, and the blue trace the beam current.

 

Linac Installation & Commissioning (Fernanda G. Garcia):

PIP2IT:

  • Main activities in PIP2IT were devoted to high beam power running conditions prior the imminent operations shutdown to start installation of CTL.
    • The ion source filament reached end of life and had to be replaced.
    • Worked on various sub-systems addressing problems as they were being revealed by increasing the pulse length.
    • By the end of the week, a beam pulse of 10msec was achieved.
  • Shutdown of PIP2IT operations has been shifted by one week or later due to Cryogenics not being ready to start the CTL installation inside the cave.
  • Internal review of the 200-ohm kicker design is tentatively scheduled for the week of May 28th.
  • Group continues working on RLS and engineering documentation.

Test Infrastructure:

  • PIP2IT Cryo transfer line installation delayed by at least one week – expect to start week of June 4th.
  • Continue working on RLS and documentation.

Installation:

  • Continued to create building-blocks for PIP-II top-level CAD model. Test case of lattice import format ongoing.

 

Conventional Facilities (Steve Dixon):

  • Site Prep Coordination Meeting – Met with the design team to coordinate progress for the Site Prep package.
  • A/E Re-compete – Tom Powers met with DOE FSO Contracting Officer to review corrections/modifications to the acquisition plan.
  • Updated conventional facilities strategy focusing on delivery of the Cryo Plant Building followed by the High Bay Building portion of the Linac Complex.